The height h (x) in centimetres of the tooh of a blade on a road-saw as it rotates x degrees is given by the function: h(x( = 60 sinx + 10
a) what is the diameter of the blade?
b) what is the maximum dept of the cut made by the blade?
c) what is the depth of the cut made by a blade (correct to the nearest tenth of a cm) when it has rotated
1) 30 degrees 2) 135 degrees 3) 240 degrees.
Thanks.

Answer by Alan3354(69443)

The height h (x) in centimetres of the tooh of a blade on a road-saw as it rotates x degrees is given by the function: h(x) = 60 sinx + 10
a) what is the diameter of the blade?
120 cm
------
b) what is the maximum dept of the cut made by the blade?
50 cm
------
c) what is the depth of the cut made by a blade (correct to the nearest tenth of a cm) when it has rotated
1) 30 degrees 2) 135 degrees 3) 240 degrees.
Angles measured from where? The direction of rotation shoule be stated, too.
You can sub the angles into the function, but it's not clear without a reference starting point.
---
Doing this implies the angles are measured from the horizontal.
h(x) = 60sin(x) + 10
h(30) = 60*0.5 + 10 = 40 cm
-------
135 & 240 degrees are past the bottom of the cut --> depth = 60 + 10 = 70 cm
======================================
This problem is not well defined. It's likely that the +10 cm is the diameter of the hub of the saw.
No reference angle is given. If the angles are those of the x-y plane, then the tooth would be in the air at 30 degs and at 135 degs.
The max depth would be 60 - 10 = 50 cm, not 70. That implies the angles are measured from the vertical, a contradiction.
=====
More info is needed.
----